# Hybrid Deepfake Image Detection: A Comprehensive Dataset-Driven Approach Integrating Convolutional and Attention Mechanisms with Frequency Domain Features

### v1

@@ 4 Conclusion
In this work, we have proposed a weighted ensemble approach that takes advantage of the strengths of a ResNet-34-based model, a DeiT model, and an XceptionNet model for the accurate classification of deepfake images. The ResNet-34 model enabled extracting local features from an image, while the DeiT model specialized in capturing global features. Additionally, XceptionNet, combined with wavelet transform, effectively identified frequency-dependent artifacts. We emphasize that detecting deepfake images that are generated with a range of generative AI approaches is a challenging task. Our proposed model architectures captured the intricate properties of the input images to successfully distinguish between real and fake images and produced excellent results. Through t-SNE plots, we demonstrated that our model is capable of separating real and fake images into separable clusters. Additionally, we included a Grad-CAM visualizations to map the focus-regions of images to illustrate how our models analyze the images and which portions contribute to distinguishing between real and fake images. An interesting future work could be the incorporation of lightweight architectures and reduce the computational overhead, as well as an increase in the model’s adaptability to newer and sophisticated deepfake generation methods.

### v2

@@ 4 Conclusion
In this work, we proposed an ensemble architecture for deepfake detection that captured the intricate properties of the input images to successfully distinguish between real and fake images and produced excellent results. We emphasize that detecting deepfake images that are generated with a range of generative AI approaches is a challenging task. To that end, we have addressed a significant class imbalance through disjoint training set-based multistage training, and proposed a weighted ensemble approach that takes advantage of the strengths of an EfficientNet model, a Data Efficient Image Transformer model, and a ConvNext model for the accurate classification of deepfake images. The EfficientNet-B0 model enabled extracting local features from an image, while the DeiT model specialized in capturing global features. Additionally, ConvNeXt, combined with the wavelet transform, effectively identified frequency-dependent artifacts. Through t-SNE plots, we demonstrated that our model is capable of separating real and fake images into separable clusters. Additionally, we included Grad-CAM visualizations to map the focus regions of images to illustrate how our models analyze the images and which portions contribute to distinguishing between real and fake images. Besides, we further analyzed the effectiveness of the proposed models by testing them against FGSM attacks and found the optimum ensemble mechanism to be majority voting compared to the weighted ensemble with a little trade-off in accuracy. An interesting future work could be incorporating lightweight architectures, reducing the computational overhead, and increasing the model’s adaptability to newer and sophisticated deepfake generation methods.